Can you please help me solve this problem? x^3-14x^2+49x+0
Answer by stanbon(57984)

x^3-14x^2+49x = 0
--------------
Factor:
x(x^2-14x+ 49) = 0
--------------
Factor again:
x(x-7)^2 = 0
x = 0 or x = 7(multiplicity two)
====================================
Cheers,
Stan H.